next five questions, numbered 121-125, require numeric answers Determine the Use the following guidelines for entering your answers he correct answer for each question and enter it in the grid on page 3 of the answer sheet Start your answer in any column, space permitting. Unused columas Start your answer in any columa. should be left blank Wrnte your answer in the boxes at the top of the grid and fill in the correspooding cirles circle in any column. You will receive credit only if the circles are filled in completely Provide your answer in the format specified by the question. The requested circles. Mark only one answer may be an integer, a decimal, or a fraction, and it may have a negative valoe. enler a fraction, use one of the division slashes to separate the numerator from the denominator, as shown in the example below. Fractions only need to be reduced enough to fit in the gndc-ain will ntr a mixed number, as this will be scored as a fraction. For example, 2 1/2 (two and one-hal?) will .Do not enter be scored as 21/2 (twenty-one halves). Decimal answer: 4.13 Fraction answer:-210 (does not have to be reduced) Integer answer: 5024 (either position is correct)Hello! Please help me with this question!! I'm desperate and have the AP test this week. A letter answer and short explanation would be so so helpful!
Answer #1

energy of producers = 25470

energy of primary consumers = 3821

now: percentage of energy tranferred to primary consumers = 3821/25470*100 = 15.00196% or 15% in round figure

energy of secondary consumers = 573

% of energy transferred to secondary consumers = 537/3821*100 = 14.99% or 15% in round figure

energy of tertiary consumers = 86

% of energy transferred to tertiary consumers = 86/573*100 = 15.0087% or 15% in round figure

so the answer is 15